# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v2

@@ 1 Introduction
Building on the Bayesian interpretation of sparse coding (olshausen1996emergence; olshausen1997sparse)—the framework that motivates SAEs—we rephrase the optimization objective of SAEs as a MAP (maximum a posteriori) estimation problem. This allows us to make explicit prior assumptions about temporal structure embedded in SAEs, showing they implicitly assume concepts are uncorrelated across time and the number of concepts necessary to explain an activation is time-invariant—that is, the information present at each token position is independent of information at other positions and uniformly distributed. As we empirically show, these assumptions stand in stark contrast to the actual temporal structure present in language and language model representations, and can result in empirically observed pathologies in SAEs, such as feature splitting (bricken2023monosemanticity; chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning). These results then motivate us to draw a broader parallel between SAEs and computational neuroscience approaches for understanding neural data. Specifically, population-level analyses of neural recordings have revealed that representations often lie on structured manifolds (khona2022attractor; nogueira2023geometry; sohn2019bayesian), challenging the reductionist assumption in sparse coding that computations occur via independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two). This motivated a paradigm shift towards more structured analysis protocols—methods designed around the generative process of the behavior one aims to explain (schneider2023learnable; chen2018sparse). Motivated by this and similar findings of intricate geometrical structure in neural network representations (fel2025into; gurnee2025when; modell2025origins), we propose Temporal Feature Analysis, a new protocol for interpreting language model activations that incorporates explicit inductive biases about temporal structure. Our approach decomposes activations at each timestep into two orthogonal components: a predictable component, obtained by projecting current representations onto past context using a learned attention mechanism, and a novel component, representing residual information orthogonal to the predictable component. That is, we assume the novel component—not the total representation—is uncorrelated over time. This allows correlations between total codes and hence enables our method to capture the temporal structure of LM activations. Overall, we make the following contributions in this paper.
@@
  • Incorporating data-informed temporal inductive biases into a novel SAE architecture (Sec. 5). Using empirically observed temporal structure in LM activations, we design a novel interpretability protocol–called Temporal Feature Analysis–which decomposes the model activation into two components: a predictive component, which captures slow-moving contextual information, and a novel component, which captures fast-moving stimulus-driven information.
@@ 3 Temporal Structure in Language Model Activations
Figure 2: Temporal structure of LLM activations reveals nonstationarity. We use Pile samples (monology2021pile-uncopyrighted) to analyze temporal structure from activations of two pretrained LMs, comparing it to a surrogate signal that is stationary in nature (see App. A.1.4). (a, e) Intrinsic dimension of model activations and stationary surrogate. (b, f) Autocorrelations $A(\bm{x}_{t},\bm{x}_{t-\tau})$ as a function of sequence position ($t$) and lag ($\tau$). (c, g) Autocorrelation of the stationary surrogate. (d, h) Variance explained by projecting current representation $\bm{x}_{t}$ onto past context window $\{\bm{x}_{t-1},\dots,\bm{x}_{t-w}\}$ with different sizes $w$, along with a baseline. Results consistently show representations getting ‘denser’ over time and being significantly more structured than a stationary surrogate.
@@
Fig. 2 (a,e) show the dimensionality of the underlying manifold structure (intrinsic dimension) in model activations. We estimate the intrinsic dimensionality at a fixed position across a set of sequences with the U-statistic (App. Sec. A.1.2). For language model activations, this metric increases steadily with sequence position. On the other hand, a stationary surrogate of the data (see App. Sec. A.1.4) shows nearly constant intrinsic dimension over time. This indicates that model activations get ‘denser’, i.e., they possess more information over time. Correspondingly, the number of concepts needed to explain them varies with context.
@@
Subplots (b), (f) show the autocorrelation of model activations (App. Sec. A.1.1), which is noticeably different at different sequence positions (x-axis), while the stationary surrogate, as expected, shows nearly position-invariant autocorrelation values (subplots (c), (g)). This finding is a clear signature of time-dependent correlation structure, and therefore of non-stationarity. We quantify the similarity of a representation with its context in subplots (d), (h). Specifically, we project representations of token $\bm{x}_{t}$ at a given time $t$ onto the subspace spanned by preceeding representations in the context $\{\bm{x}_{<t}\}$. These subplots show that up to $80\%$ variance of $\bm{x}_{t}$ is explained by a context of 500 tokens, further highlighting strong cross-temporal correlations. Significant variance in the representation at time $t$, $\bm{x}_{t}$, can be predicted (expressed) using representations from the past context.
@@ 4 Temporal Priors of Sparse Autoencoders
A more precise version of the claim for specific SAE architectures is provided in Appendix G.1. Intuitively, the claim above says that SAEs assume an independence of latents, and hence the concepts underlying the generative process of language, over time. This directly conflicts with the rich contextual structure of LM activations we empirically observed in Fig. 2b–d, f–h. Crucially, this also implies that SAEs assume the sparsity of latent codes necessary to explain model activations to be time-invariant, as stated formally in the corollary below.
@@
Intuitively, the claim above says that even if an SAE achieves low reconstruction error, its specific geometric assumptions—use of independently firing directions and restricted intrinsic dimensionality of data—will lead to a “splitting” of the input data into several regions such that different sets of latents will activate for even locally nearby points. Consequently, if the sparsity budget is too little or if model activations are more dense than the assumed budget, then local geometry present in the activations will be lost in the SAE latent codes—as is arguably already observed empirically with phenomena like feature splitting (chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning; bricken2023monosemanticity). For the context of our paper, as per our results in Fig. 2 (a,e), temporal correlations can increase activations’ intrinsic dimensionality and, when viewed from the lens of the claim above, we arrive at the conclusion that SAEs can fail to capture temporal structure present in activations. We empirically validate this claim in Sec. 6. More broadly, we note there is growing evidence that neural network representations possess rich geometric structures across a broad set of modalities (gurnee2025when; fel2025into; park2025iclrincontextlearningrepresentations; engels2024not; costa2025flat; kantamneni2025language; pearce2025tree). As per Prop. 4.2, the use of SAEs for interpreting behaviors that invoke such structures may yield misleading explanations and should be approached carefully.
@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
As stated in Sec. 2, sparse coding, a framework designed in computational neuroscience to understand neural representations in biological brains (olshausen1996emergence; olshausen1997sparse), inspired SAEs as a framework for interpreting artificial neural networks (bricken2023monosemanticity; cunningham2023sparse). In fact, the parallels between these communities can be made deeper: motivated by observations of intricate geometry of neural representations derived out of multi-dimensional population analyses (khona2022attractor; nogueira2023geometry; sohn2019bayesian), there were calls in computational neuroscience to discard the limiting reduction assumed in sparse coding that computations occur via a set of independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two; seung1996brain; chung2021neural)—similar to our arguments in Sec. 34 (and results that follow in Sec. 6). Correspondingly, a need for more structured protocols was suggested (eichenbaum2018barlow; barack2021two; chen2019sparse), leading to methods that were motivated by the generative process of the behavior one is trying to explain (schneider2023learnable; chen2018sparse; berkes2005slow; wiskott2002slow; linderman2017bayesian; yu2008gaussian). We argue a similar paradigm shift is needed in language model interpretability: given that we train models to learn the distribution of highly structured data, we ought to embrace the fact that neural network activations can exhibit intricate geometrical organization. These geometrical objects may in fact be the units of computation that are necessary for describing model computation (perich2025neural; eichenbaum2018barlow; modell2025origins), and hence we must design interpretability approaches to identify them. In what follows, as an attempt to qualify our arguments, we propose one such approach, titled Temporal Feature Analysis, that focuses on the temporal structure of LM activations.
@@
In computational neuroscience, when analyzing data from dynamical domains (e.g., audio, language, or video), a commonly made assumption is that there is contextual information present in the recent history that informs the next state—this part of the signal is deemed predictable (chen2025context; millidge2024predictive), slow-changing (berkes2005slow), invariant (olshausen2007learning), or dense (tasissa2022improvingdiscriminativereconstructionsimultaneous). Meanwhile, the remaining signal corresponds to new bits of information added by the observed state at the next timestep—this part can be deemed novel, fast-changing, variant, or sparse with respect to the context. We argue LM activations are amenable to a similar generative model. Specifically, our observations in Sec. 3 show that activations $\bm{x}_{t}$ at time $t$ are strongly correlated with the context and can be decomposed into two such parts as well. Motivated by this, we propose the following model of LM activations:
$\bm{x}_{t}=\bm{x}_{p,t}+\bm{x}_{n,t},\quad\text{s.t. }\,\,\bm{x}_{n,t}\perp\bm{x}_{p,t},\,\,\bm{x}_{p,t}=\bm{D}\bm{z}_{p,t},\,\,\text{and}\,\,\bm{x}_{n,t}=\bm{D}\bm{z}_{n,t},$ (3)
@@
where $\bm{x}_{p,t}=\bm{D}\bm{z}_{p,t}$ denotes a predictable component of the signal that captures the correlations of $\bm{x}_{t}$ with past data $\left\{\bm{x}_{<t}\right\}$, while $\bm{x}_{n,t}=\bm{D}\bm{z}_{n,t}$ denotes a novel component that represents new information added by the current token $\bm{x}_{t}$. To obtain $\bm{z}_{p,t}$, we project $\bm{x}_{t}$ onto $\{\bm{x}_{<t}\}$ to explain the predictable variance in $\bm{x}_{t}$ as a convex combination of past data. Specifically, we use a self-attention layer $f$ on top of a single ReLU layer, yielding $\bm{z}_{p,t}=f(\ \{\bm{x}_{1},\dots,\bm{x}_{t-1}\},\bm{x}_{t})$. Meanwhile, $\bm{z}_{n,t}=\tilde{f}(\bm{x}_{t},\bm{z}_{p,t})=\sigma(\bm{D}^{T}(\bm{x}_{t}-\bm{D}\bm{z}_{p,t}))$ captures the residual component of the code, which is not correlated with the past (note that $\sigma$ is the nonlinearity). We use a standard SAE encoder (either TopK or BatchTopK) to instantiate $\sigma$, applying it to $\bm{x}_{t}-\bm{D}\bm{z}_{p,t}$ to derive $\bm{z}_{n,t}$. See Fig. 4 for an overall schematic of the encoding process. The learning objective in Temporal Feature Analysis follows.
@@
$\begin{split}\operatorname*{arg\,min}_{\bm{D},\bm{z}}\frac{1}{N}&\sum_{i=1}^{T}\|\bm{x}_{i}-\bm{D}(\bm{z}_{p,i}+\bm{z}_{n,i})\|_{2}^{2}+\lambda\mathcal{R}(\bm{z}_{n,i}),\\ \text{s.t. }\bm{z}_{p,k}=f_{\mathtt{SAE}}(\{\bm{x}_{<k}\},\bm{x}_{k})&,\;\bm{z}_{n,k}=\tilde{f}_{\mathtt{SAE}}(\bm{x}_{k},\bm{z}_{p,k}),\;\bm{z}_{k}=\bm{z}_{p,k}+\bm{z}_{n,k}\;\forall k.\end{split}$ (4)
@@
Relating to Sec. 4, we note the prior assumption in Temporal Feature Analysis is that the residual $\bm{z}_{n,t}=\bm{z}_{t}-\bm{z}_{p,t}$, which captures the novel information in $\bm{x}_{t}$ remaining after removing the projections onto the past context, is i.i.d. over time. This prior allows correlations between total codes $\bm{z}$ at different times, and thereby allows correlations between concepts across time, instead of assuming that all concepts are time-independent. We refer the reader to Sec. 8 for a discussion on limitations of this assumed generative model.
@@
Table 3: predictive and novel codes explain different parts of the input signal across domains.
Sim. % Norm NMSE Var. Expl.

### v3

@@ 1 Introduction
Building on the Bayesian interpretation of sparse coding (olshausen1996emergence; olshausen1997sparse)—the framework that motivates SAEs—we rephrase the optimization objective of SAEs as a MAP (maximum a posteriori) estimation problem. This allows us to make explicit prior assumptions about temporal structure embedded in SAEs, showing they implicitly assume concepts are uncorrelated across time and the number of concepts necessary to explain an activation is time-invariant—that is, the information present at each token position is independent of information at other positions and uniformly distributed. As we empirically show, these assumptions stand in stark contrast to the actual temporal structure present in language and language model representations, and can result in empirically observed pathologies in SAEs, such as feature splitting (bricken2023monosemanticity; chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning).
These results then motivate us to draw a broader parallel between SAEs and computational neuroscience approaches for understanding neural data. Specifically, population-level analyses of neural recordings have revealed that representations often lie on structured manifolds (khona2022attractor; nogueira2023geometry; sohn2019bayesian), challenging the reductionist assumption in sparse coding that computations occur via independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two). This motivated a paradigm shift towards more structured analysis protocols—methods designed around the generative process of the behavior one aims to explain (schneider2023learnable; chen2018sparse). Motivated by this and similar findings of intricate geometrical structure in neural network representations (fel2025into; gurnee2025when; modell2025origins), we propose Temporal Feature Analysis, a new protocol for interpreting language model activations that incorporates explicit inductive biases about temporal structure. Our approach decomposes activations at each timestep into two orthogonal components: a predictable component, obtained by projecting current representations onto past context using a learned attention mechanism, and a novel component, representing residual information orthogonal to the predictable component. That is, we assume the novel component—not the total representation—is uncorrelated over time. This allows correlations between total codes and hence enables our method to capture the temporal structure of LM activations. Overall, we make the following contributions in this paper.
@@
  • Incorporating data-informed temporal inductive biases into a novel SAE architecture (Sec. 5). Using empirically observed temporal structure in LM activations, we design a novel interpretability protocol—called Temporal Feature Analysis—which decomposes the model activation into two components: a predictive component, which captures slow-moving contextual information, and a novel component, which captures fast-moving stimulus-driven information.
@@ 3 Temporal Structure in Language Model Activations
Figure 2: Temporal structure of LLM activations reveals nonstationarity. We use Pile samples (monology2021pile-uncopyrighted) to analyze temporal structure from activations of two pretrained LMs, comparing it to a surrogate signal that is stationary in nature (see App. A.4). (a, e) Intrinsic dimension of model activations and stationary surrogate. (b, f) Autocorrelations $A(\bm{x}_{t},\bm{x}_{t-\tau})$ as a function of sequence position ($t$) and lag ($\tau$). (c, g) Autocorrelation of the stationary surrogate. (d, h) Variance explained by projecting current representation $\bm{x}_{t}$ onto past context window $\{\bm{x}_{t-1},\dots,\bm{x}_{t-w}\}$ with different sizes $w$, along with a baseline. Results consistently show representations getting ‘denser’ over time and being significantly more structured than a stationary surrogate.
@@
Fig. 2 (a,e) show the dimensionality of the underlying manifold structure (intrinsic dimension) in model activations. We estimate the intrinsic dimensionality at a fixed position across a set of sequences with the U-statistic (App. Sec. A.3). For language model activations, this metric increases steadily with sequence position. On the other hand, a stationary surrogate of the data (see App. Sec. A.4) shows nearly constant intrinsic dimension over time. This indicates that model activations get ‘denser’, i.e., they possess more information over time. Correspondingly, the number of concepts needed to explain them varies with context.
@@
Subplots (b), (f) show the autocorrelation of model activations (App. Sec. A.1), which is noticeably different at different sequence positions (x-axis), while the stationary surrogate, as expected, shows nearly position-invariant autocorrelation values (subplots (c), (g)). This finding is a clear signature of time-dependent correlation structure, and therefore of non-stationarity. We quantify the similarity of a representation with its context in subplots (d), (h). Specifically, we project representations of token $\bm{x}_{t}$ at a given time $t$ onto the subspace spanned by preceeding representations in the context $\{\bm{x}_{<t}\}$. These subplots show that up to $80\%$ variance of $\bm{x}_{t}$ is explained by a context of 500 tokens, further highlighting strong cross-temporal correlations. Significant variance in the representation at time $t$, $\bm{x}_{t}$, can be predicted (expressed) using representations from the past context.
@@ 4 Temporal Priors of Sparse Autoencoders
A more precise version of the claim for specific SAE architectures is provided in Appendix H.1. Intuitively, the claim above says that SAEs assume an independence of latents, and hence the concepts underlying the generative process of language, over time. This directly conflicts with the rich contextual structure of LM activations we empirically observed in Fig. 2b–d, f–h. Crucially, this also implies that SAEs assume the sparsity of latent codes necessary to explain model activations to be time-invariant, as stated formally in the corollary below.
@@
Intuitively, the claim above says that even if an SAE achieves low reconstruction error, its specific geometric assumptions—use of independently firing directions and restricted intrinsic dimensionality of data—will lead to a “splitting” of the input data into several regions such that different sets of latents will activate for even locally nearby points (see Fig. 3). Consequently, if the sparsity budget is too little or if model activations are more dense than the assumed budget, then local geometry present in the activations will be lost in the SAE latent codes—as is arguably already observed empirically with phenomena like feature splitting (chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning; bricken2023monosemanticity). For the context of our paper, as per our results in Fig. 2 (a,e), temporal correlations can increase activations’ intrinsic dimensionality and, when viewed from the lens of the claim above, we arrive at the conclusion that SAEs can fail to capture temporal structure present in activations. We empirically validate this claim in Sec. 6. More broadly, we note there is growing evidence that neural network representations possess rich geometric structures across a broad set of modalities (gurnee2025when; fel2025into; park2025iclrincontextlearningrepresentations; engels2024not; costa2025flat; kantamneni2025language; pearce2025tree). As per Prop. 4.2, the use of SAEs for interpreting behaviors that invoke such structures can be misleading and should be approached carefully (cf. (sengupta2018manifold)).
@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
As stated in Sec. 2, sparse coding, a framework designed in computational neuroscience to understand neural representations in biological brains (olshausen1996emergence; olshausen1997sparse), inspired SAEs as a framework for interpreting artificial neural networks (bricken2023monosemanticity; cunningham2023sparse). In fact, the parallels between these communities can be made deeper: motivated by observations of intricate geometry of neural representations derived out of multi-dimensional population analyses (khona2022attractor; nogueira2023geometry; sohn2019bayesian), there were calls in computational neuroscience to discard the limiting reduction assumed in sparse coding that computations occur via a set of independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two; seung1996brain; chung2021neural)—similar to our arguments in Sec. 34 (and results that follow in Sec. 6). Correspondingly, a need for more structured protocols was suggested (eichenbaum2018barlow; barack2021two; chen2019sparse; sengupta2018manifold), leading to methods that were motivated by the generative process of the behavior one is trying to explain (schneider2023learnable; chen2018sparse; berkes2005slow; wiskott2002slow; linderman2017bayesian; yu2008gaussian). We argue a similar paradigm shift is needed in language model interpretability: given that we train models to learn the distribution of highly structured data, we ought to embrace the fact that neural network activations can exhibit intricate geometrical organization. These geometrical objects may in fact be the units of computation that are necessary for describing model computation (perich2025neural; eichenbaum2018barlow; modell2025origins), and hence we must design interpretability approaches to identify them. In what follows, as an attempt to qualify our arguments, we propose one such approach, titled Temporal Feature Analysis, that focuses on the temporal structure of LM activations. We also point the reader to relevant related work from identifiability literature (joshi2025identifiable; klindt2020towards).
@@
In computational neuroscience, when analyzing data from dynamical domains (e.g., audio, language, or video), a commonly made assumption is that there is contextual information present in the recent history that informs the next state—this part of the signal is deemed predictable (chen2019sparse; millidge2024predictive), slow-changing (berkes2005slow; klindt2020towards), invariant (olshausen2007learning; hyvarinen2003bubbles), or dense (tasissa2022improvingdiscriminativereconstructionsimultaneous). Meanwhile, the remaining signal corresponds to new bits of information added by the observed state at the next timestep—this part can be deemed novel, fast-changing, variant, or sparse with respect to the context. We argue LM activations are amenable to a similar generative model. Specifically, our observations in Sec. 3 show that activations $\bm{x}_{t}$ at time $t$ are correlated with the context and can be decomposed into two such parts as well. Motivated by this, we propose the following model of LM activations:
$\bm{x}_{t}=\bm{x}_{p,t}+\bm{x}_{n,t},\quad\text{s.t. }\quad\bm{x}_{p,t}=\bm{D}\bm{z}_{p,t}\,\,\text{and}\,\,\bm{x}_{n,t}=\bm{D}\bm{z}_{n,t},$ (3)
@@
where $\bm{x}_{p,t}$ denotes a predictable component of the signal that captures the correlations of $\bm{x}_{t}$ with past data $\left\{\bm{x}_{<t}\right\}$, while $\bm{x}_{n,t}$ denotes a novel component that represents new, i.e., uncorrelated with the past, information added by the current token $\bm{x}_{t}$. To obtain $\bm{z}_{p,t}$, we project $\bm{x}_{t}$ onto $\{\bm{x}_{<t}\}$ to explain the predictable variance in $\bm{x}_{t}$ as a convex combination of past data. Specifically, we apply an encoder (ReLU plus a linear map) to take the inputs to a latent space $\bm{v}$ and perform an attention operation $f$ in that space, yielding $\bm{z}_{p,t}=f(\ \{\phi(D^{T}\bm{x}_{1}),\dots,\phi(D^{T}\bm{x}_{t-1})\},\phi(D^{T}\bm{x}_{t}))$; here $\bm{x}_{t}$ defines the query and the remaining context serves as keys. Meanwhile, $\bm{z}_{n,t}$ is defined using a standard SAE: $\bm{z}_{n,t}=\tilde{f}_{\mathtt{SAE}}(\bm{x}_{t},\bm{z}_{p,t})=\sigma(\bm{D}^{T}(\bm{x}_{t}-\bm{D}\bm{z}_{p,t}))$; we use a standard SAE nonlinearity (either TopK or BatchTopK) to instantiate $\sigma$, applying it to $\bm{x}_{t}-\bm{D}\bm{z}_{p,t}$ to derive $\bm{z}_{n,t}$. See Fig. 4 for an overall schematic of the encoding process. The learning objective in Temporal Feature Analysis follows.
@@
$\begin{split}\operatorname*{arg\,min}_{\bm{D},\bm{z}}\frac{1}{T}&\sum_{i=1}^{T}\|\bm{x}_{i}-\bm{D}(\bm{z}_{p,i}+\bm{z}_{n,i})\|_{2}^{2}+\lambda\mathcal{R}(\bm{z}_{n,i}),\\ \text{s.t. }\,\,\,\bm{z}_{p,k}&=f_{\mathtt{SAE}}(\{\bm{x}_{<k}\}\bm{x}_{k})\,\,\,\text{and}\,\,\bm{z}_{n,k}=\tilde{f}_{\mathtt{SAE}}(\bm{x}_{k},\bm{z}_{p,k})\,\,\forall k.\end{split}$ (4)
@@
Relating to Sec. 4, we note the prior assumption in Temporal Feature Analysis is that the residual $\bm{z}_{n,t}=\bm{z}_{t}-\bm{z}_{p,t}$, which captures the novel information in $\bm{x}_{t}$ remaining after removing the projections onto the past context, is i.i.d. over time. This prior allows correlations between the total codes, and thereby between concepts, across time, instead of assuming that all concepts are time-independent. However, we do note this prior is still an assumption: one can reasonably argue the residuals need not be i.i.d.. We nevertheless state this point to be explicit about assumptions involved in our work.
@@
Table 3: Predictive and novel codes explain different parts of the input. See main text for details.
$\langle\bm{x}_{p},\bm{x}_{n}\rangle$ % Norm NMSE Var. Expl.